OpenStudy (anonymous):

What is 3/5 divided by 1/2

OpenStudy (raffle_snaffle):



OpenStudy (anonymous):

6 / 5

OpenStudy (anonymous):

6/5

OpenStudy (jdoe0001):

\(\bf \cfrac{\frac{3}{5}}{\frac{1}{2}}\implies \cfrac{3}{5}\times \cfrac{2}{1}\\ \quad \\ \textit{recall that}\quad \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\)